OpenStudy (anonymous):

the ratio of men to women in class is 6 to 5. how many women students are there if there are 3600 men? please show work

OpenStudy (anonymous):

this measn that there are \(\frac{6}{5}\) times more men then women, so if you do \[\frac{3600}{\frac{6}{5}} = females = 3000\]

OpenStudy (anonymous):

ok thank u